OpenStudy (apex128):

Guys i really need help w/ this question! A 203.7-kg satellite is orbiting planet Y at an altitude of 1.3x10^6 meters above its surface and with an orbital period of 19.4 hours. If the planet has a mass of 1.9x10^24 kg, then determine the radius of planet Y.

OpenStudy (steve816):

(GMm)/(r^2)=mg

OpenStudy (apex128):

thx 8

OpenStudy (apex128):

im going to fail this test btw.

OpenStudy (steve816):

Well, you won't if you study.

OpenStudy (apex128):

It says it wrong

OpenStudy (apex128):

what do i put for little g

OpenStudy (steve816):

Nevermind sorry, you had to use this equation: (GM)/(4pi^2) = (R^3)(T^2).